Am I asking this correctly? The directions say: Use the substitution method to solve the linear system.
Problem:
1. y = x + 2
2x + y = 8

Answer by stanbon(57333)

y = x + 2
2x + y = 8
-------------
Solve both for "y":
y = x+2
y = -2x+8
-------------
Substitute "x+2" in place of "y" in the 2nd equation:
x+2 = -2x+8
3x = 6
x=2
---------
Substitute into y = x+2 to solve for y:
y = 2+2 = 4
==============================
Cheers,
Stan H.